# The Overcooked Generalisation Challenge

### v2

@@ Abstract
We introduce the Overcooked Generalisation Challenge (OGC) – the first benchmark to study reinforcement learning agents’ zero-shot cooperation abilities when faced with novel partners and levels in the Overcooked-AI environment. This perspective starkly contrasts a large body of previous work that evaluated cooperating agents only on the same level or with the same partner, thus failing to capture generalisation abilities essential for real-world human-AI cooperation. Our challenge interfaces with state-of-the-art dual curriculum design (DCD) methods to generate auto-curricula for training general agents in Overcooked. OGC is the first environment specially designed for DCD methods and, consequently, the first evaluated with state-of-the-art methods. It is fully GPU-accelerated, built on the DCD benchmark suite minimax, and freely available as open source. We show that state-of-the-art DCD algorithms fail to produce useful policies on OGC, even if combined with recent network architectures specifically designed for scalability and generalisability. As such, the OGC pushes the boundaries of real-world human-AI cooperation by enabling pioneering research on the impact of generalisation on cooperating agents.
Human-AI Cooperation, Multi-Agent Reinforcement Learning, Unsupervised Environment Design, Benchmark
@@ 1 Introduction
Figure 1: The Overcooked Generalisation Challenge (OGC) is the first benchmark that requires reinforcement learning agents to cooperate with novel partners in novel levels in a zero-shot fashion. During training, agents can access a generator that outputs new training environments. During evaluation, agents are presented with a novel level and partner to cooperate with.
Developing computational agents capable of collaborating with humans has emerged as a key challenge in artificial intelligence (AI) research (Stone et al., 2010; Dafoe et al., 2020) and promises to vastly expand human abilities (O’neill et al., 2020). Recent years have seen considerable advances in understanding human cooperative behaviour (Rand & Nowak, 2013; Vizmathy et al., 2024), computational modelling of cooperation (Nikolaidis & Shah, 2013; Sadigh et al., 2016; Ding et al., 2024), as well as in developing computational methods for human-AI cooperation (Hu et al., 2020; Strouse et al., 2021). In parallel, several benchmarks were proposed to foster the development and evaluation of these methods (Samvelyan et al., 2019; Bard et al., 2020). Most notably, Overcooked-AI (Carroll et al., 2019) has established itself as a widely used benchmark for evaluating (zero-shot) human-AI coordination (Strouse et al., 2021; Zhao et al., 2023; Yu et al., 2023).
Despite the advances they have enabled, all of these benchmarks are limited in that they only allow to assess reinforcement learning (RL) agents’ cooperative abilities in-distribution. That is, they either only allow to evaluate agents in the same environment in which they were trained (Hu et al., 2020; Carroll et al., 2019) or the same partner agent they were trained with (Foerster et al., 2018; Lowe et al., 2017; Strouse et al., 2021). In Overcooked-AI, for instance, existing zero-shot coordination (ZSC) methods are trained once per layout at considerable computational cost (Carroll et al., 2019; Yang et al., 2022; Zhao et al., 2023; Yu et al., 2023), and these layouts only feature a limited number of possible cooperation strategies. However, real collaborative settings require coordination with novel partners and in unknown environments. While previous open-source benchmarks studied opponents combined with map generalisation, no dedicated open-source benchmarks exist for studying cooperation partners combined with map generalisation. Consider a medical robot assisting doctors in hospitals as an example of why this might be useful. Such a robot will be deployed in unique and unknown hospitals and surgical rooms, where it must adapt to different medical staff and their preferences.
To address this limitation, we introduce the Overcooked Generalisation Challenge (OGC) – the first zero-shot cooperation benchmark that challenges agents to cooperate in novel layouts and with unknown partners, see Figure 1. Cooperative settings differ from competitive ones in their game-theoretic background and thus require separate algorithms and benchmarks (Lerer & Peysakhovich, 2019). Neural MMO (Suarez et al., 2021, 2023) comes closest as it mixes cooperation and competition but – unlike Overcooked-AI – does not provide a purely cooperative setting that is specifically designed for human-AI collaboration.
To train and evaluate agents on OGC, we make use of unsupervised environment design (UED) (Dennis et al., 2020) to generate suitable training levels, provide hand-designed testing levels, assess ad-hoc teamwork on these by providing populations of diverse testing agents and then similarly also test for zero-shot cooperation. As such, our work is the first to combine UED techniques with a multi-agent RL zero-shot cooperation task and thus bridges the gap between two previously unrelated research areas; it studies the impact of generalisation on human-AI coordination and the ability of UED algorithms to design optimal auto-curricula for cooperating agents. We show that several state-of-the-art UED algorithms and network architectures struggle to perform well on our challenge. Only PAIRED (Dennis et al., 2020) together with a policy that incorporates a soft Mixture-of-Experts (SoftMoE) module (Obando-Ceron et al., 2024) has limited success at generalising to the testing levels and outperforms competitive baselines, including robust PLR (Jiang et al., 2021b, a) and ACCEL (Parker-Holder et al., 2022). Our findings call for future work to develop new methods combining zero-shot coordination and DCD techniques in a single ZSC-DCD framework. Our benchmark provides the necessary environment to facilitate these developments. Taken together, our contributions are three-fold:
@@
    We introduce the Overcooked Generalisation Challenge – a novel benchmark challenge in which agents are asked to cooperate with novel partners and in previously unseen layouts.
@@
    We provide OvercookedUED – an open-source environment that can be used with state-of-the-art DCD algorithms and that is integrated into minimax (Jiang et al., 2023), taking full advantage of hardware acceleration provided by JAX.
@@
    We establish baseline performance on the OGC using state-of-the-art DCD algorithms and neural network architectures (Smith et al., 2023; Obando-Ceron et al., 2024). We show that current DCD algorithms struggle with our new challenge, including the most recent network architectures. Notably, we show the generalisation performance of SoftMoE-based architectures in hard generalisation tasks. Furthermore, we assess zero-shot cooperation performance with a population of diverse partners to link zero-shot cooperation and generalisation. We show that, as policies become more generally capable, they achieve better zero-shot cooperation.
2 Research Challenges
Figure 2: The OGC features a challenging large design space where many different elements have to be placed in relation to each other. This makes the OGC challenging for both designing learning levels and training agents.
The OGC poses several new challenges for zero-shot human-AI cooperation that go beyond existing benchmarks and that are essential for further advances in the development of cooperating RL agents:
Generalisation
The OGC challenges the generalisation capabilities of methods and agents by having them engage in a double generalisation challenge: adjusting to both novel partners and levels. Existing cooperative open-source benchmarks typically require only one form of generalisation; see, for instance, (Lowe et al., 2017; Foerster et al., 2018; Carroll et al., 2019; Hu et al., 2020).
Environment Design
Our environment challenges UED algorithms in generating and designing layouts with many interacting components and agents. This is in contrast to existing environments that only require UED algorithms to design simple mazes, 2D walker terrains, or race tracks consisting of fewer elements (Dennis et al., 2020; Jiang et al., 2021a; Parker-Holder et al., 2022; Rutherford et al., 2024a). We show that methods struggle to design layouts similar to the ones humans designed. Current methods specifically fail to design layouts requiring handing over items over a countertop or featuring deliberately designed circuits. Our benchmark challenges further research to develop UED methods that design more realistic collaboration environments for curriculum learning, possibly along the lines of Bruce et al. (2024).
Combining Environment and Partner Generalisation
Coordinating with novel partners and generalising to novel levels were often treated as separate tasks. Previous work on methods for zero-shot coordination often requires training a diverse population of partners. Since there currently is no algorithm to train a population of diverse agents over a distribution of levels, these methods need adjusting to be applicable to our challenge. Our benchmark thus encourages merging both lines of research and developing UED-ZSC methods, i.e. methods that learn both at the same time, in a single joint generalisation task.
@@ 3 Related work
3.1 Generalisation in Reinforcement Learning
@@
A large number of works have shown that RL agents fail to generalise to new environments (Zhang et al., 2018a; Cobbe et al., 2019) and have triggered research on the generalisation capabilities of RL agents (Nichol et al., 2018; Cobbe et al., 2019, 2020). Early results revealed that RL agents can memorise large numbers of levels during training (Zhang et al., 2018b; Cobbe et al., 2019) and that they must experience sufficiently diverse training data to generalise well (Cobbe et al., 2020). One established approach to generate diverse training data is to use domain randomisation (Jakobi, 1997, DR). Still, DR has been shown to produce many uninformative samples (Khirodkar et al., 2018), which can lead to the agent’s inability to generalise (Dennis et al., 2020).
3.2 Unsupervised Environment Design
Intending to address this challenge, later works on generalisation focused on unsupervised environment design (Dennis et al., 2020, UED). UED aims to improve domain randomisation by generating auto-curricula that include training levels of increasing complexity to facilitate continued agent learning (Graves et al., 2017). It does so by adapting the free parameters of an under-specified environment to the agent’s capabilities. Most popular UED methods fall into the category of Dual Curriculum Design (Jiang et al., 2021a, DCD) that combine 1) an agent, 2) a level generator, and 3) a curator that picks which levels to train on. Popular methods include Prioritised Level Replay (PLR) (Jiang et al., 2021b), robust PLR (Jiang et al., 2021a), MAESTRO (Samvelyan et al., 2023), ReMiDi (Beukman et al., 2024), PAIRED (Dennis et al., 2020), ACCEL (Parker-Holder et al., 2022), and Replay-Enhanced (RE)PAIRED (Jiang et al., 2021a). While the development of these DCD methods has been steady, they have mostly been explored in simple single-agent environments For a comprehensive overview, see Table 3 and Appendix A.3. Compared to single-agent environments, multi-agent environments are inherently more complex because the agents interact with each other, as well as with the environment. Multi-agent environments are still rarely used in UED research. Most prominent is Deepmind’s XLand (Team et al., 2021; Bauer et al., 2023), a closed-source multi-task universe for generating single- and multi-agent tasks and environments. While XLand features cooperative tasks, it is not available to researchers for studying cooperative multi-agent UED algorithms. While an open-source variant was recently published (Nikulin et al., 2023), it only supports a single agent. Arguably closest is Neural MMO (Suarez et al., 2021, 2023), a massively multi-task and multi-agent environment that combines cooperation and competition to replicate massively multiplayer online games. Instead, we are interested in assessing and identifying cooperation performance in specially designed human-AI cooperation challenges for which the massive multi-tasking and multi-agent cooperation-competition setting of Neural MMO is unsuitable. Additionally, classic Overcooked benefits from a rich history of human-AI cooperation research. Finally, while JaxNav (Rutherford et al., 2024a) features multi-agent path-finding, no interaction between agents is required, and the environment is not focused on human-AI cooperation. Other open-source environments are competitive, i.e. LaserTag (Lanctot et al., 2017; Samvelyan et al., 2023) and MultiCarRacing (Schwarting et al., 2021; Samvelyan et al., 2023), and thus not applicable to our setting. Opposed to all of these, our work contributes to and analyses the first open-source cooperative multi-agent UED environment.
3.3 Human-AI Cooperation in Overcooked-AI
Overcooked-AI (Carroll et al., 2019) has become one of the most important benchmarks for human-AI cooperation. The environment is fully cooperative and has two agents cook and deliver soups to earn a joint reward. Overcooked-AI was, for example, used in research on zero-shot cooperation (Strouse et al., 2021; Zhao et al., 2023; Yu et al., 2023; Li et al., 2023b; Yan et al., 2023, ZSC), language model-based cooperative agents (Liu et al., 2024; Tan et al., 2024), human modelling in cooperation (Yang et al., 2022). Zero-shot cooperation (Hu et al., 2020) refers to cooperating with a partner not encountered during training. It is an important proxy to ensure the ability of an agent to coordinate with humans at test time, given that human data is often unavailable and agents thus must be able to coordinate effectively without previous training. It is commonly studied in Overcooked. Our work is thus the first to explore the impact of cross-level generalisation for zero-shot cooperation and is the first to provide the necessary tools for this.
@@ 4 Preliminaries
The cooperative multi-agent UED setting can be formalised as a decentralised under-specified partially observable Markov decision process (Dec-UPOMDP) with shared rewards. A Dec-UPOMDP is defined as $\mathcal{M}=\langle\mathcal{N},A,\Omega,\Theta,\mathcal{S}^{\mathcal{M}},\mathcal{T}^{\mathcal{M}},O^{\mathcal{M}},\mathcal{R}^{\mathcal{M}},\gamma\rangle$ in which $\mathcal{N}$ is the set of agents with cardinality $n$, $\Omega$ is a set of observations, and $\mathcal{S}^{\mathcal{M}}$ is the set of true states in the environment. Partial observations $o^{i}\in\Omega$ are obtained by agent $i\in\mathcal{N}$ using the observation function $O:\mathcal{S}\times\mathcal{N}\rightarrow\Omega$. Following (Jiang et al., 2021a), a level $\mathcal{M}_{\theta}$ is defined as a fully-specified environment given some parameters $\theta\in\Theta$. In it, agents each pick an action $a_{i}\in A$ simultaneously to produce a joint action $\boldsymbol{a}=(a_{1},\dots,a_{n})$ and observe a shared immediate reward $R(s,\boldsymbol{a})$. Then, the environment transitions to the next state according to a transition function $\mathcal{T}:\mathcal{S}\times\mathcal{A}^{1}\times...\times\mathcal{A}^{n}\times\Theta\rightarrow\Delta(\mathcal{S})$ where $\Delta(\mathcal{S})$ refers to the space of distributions over $\mathcal{S}$. $\gamma\in[0,1)$ specifies the discount factor. Agents learn a policy $\pi$. The joint policy $\boldsymbol{\pi}$ together with the discounted return $R_{t}=\sum^{\infty}_{i=0}\gamma^{i}r_{t+1}$ induce a joint action value function $Q^{\boldsymbol{\pi}}=\mathbb{E}_{s_{t+1:\infty},\boldsymbol{a}_{t+1:\infty}}[R_{t}|s_{t},\boldsymbol{a}_{t}]$. The definition of the Dec-UPOMDP extends a Dec-POMDP (Oliehoek & Amato, 2016; Wu et al., 2021) with the free parameters of the environment $\Theta$, analogously to previous works (Dennis et al., 2020; Jiang et al., 2021a; Samvelyan et al., 2023).
Our definition differs from (Samvelyan et al., 2023) in terms of the shared rewards and general-sum nature. Within our Dec-UPOMDP, we perform UED to train a policy over a distribution of fully specified environments that enable optimal learning. This is facilitated by obtaining an environment policy $\Lambda$ (Dennis et al., 2020) that specifies a sequence of environment parameters $\Theta^{T}$ for the given policy that is to be trained. How $\Lambda$ is obtained depends on the DCD method. In OvercookedUED, $\Theta$ represents the possible positions of walls, pots, serving spots, agent starting locations, and onion and bowl piles adjusted by $\Lambda$ throughout training.

### v3

@@ Abstract
We introduce the Overcooked Generalisation Challenge (OGC) – a new benchmark for evaluating reinforcement learning (RL) agents on their ability to cooperate with unknown partners in unfamiliar environments. Existing work typically evaluated cooperative RL only in their training environment or with their training partners, thus seriously limiting our ability to understand agents’ generalisation capacity – an essential requirement for future collaboration with humans. The OGC extends Overcooked-AI to support dual curriculum design (DCD). It is fully GPU-accelerated, open-source, and integrated into the minimax DCD benchmark suite. Compared to prior DCD benchmarks, where designers manipulate only minimal elements of the environment, OGC introduces a significantly richer design space: full kitchen layouts with multiple objects that require the designer to account for interaction dynamics between agents. We evaluate state-of-the-art DCD algorithms alongside scalable neural architectures and find that current methods fail to produce agents that generalise effectively to novel layouts and unfamiliar partners. Our results indicate that both agents and curriculum designers struggle with the joint challenge of partner and environment generalisation. These findings establish OGC as a demanding testbed for cooperative generalisation and highlight key directions for future research. We open-source our code [^1] Our project web-page is accessible at https://collaborative-ai.org/publications/ruhdorfer25_tmlr/..
@@ 1 Introduction
Developing computational agents capable of collaborating with each other has emerged as a key challenge in artificial intelligence (AI) research (Dafoe et al., 2020). Recent years have seen considerable advances in developing cooperative reinforcement learning (RL) agents (Stone et al., 2010; Hu et al., 2020; Choudhury et al., 2020; Ding et al., 2024) and several benchmarks were proposed to evaluate their generalisation abilities (Samvelyan et al., 2019; Bard et al., 2020). However, these benchmarks typically treat generalisation to novel environments (Cobbe et al., 2019) and novel partners (Hu et al., 2020; Carroll et al., 2019) as distinct challenges. Yet, future human-AI collaboration will require agents to generalise along both axes simultaneously. For instance, an autonomous robot assisting in a disaster response team must coordinate with ever-changing human partners in unfamiliar, dynamic environments.
Overcooked-AI (Carroll et al., 2019) has emerged as one of the most popular benchmarks for evaluating zero-shot coordination. Nonetheless, agents are typically trained and evaluated on a few fixed layouts (Strouse et al., 2021; Yang et al., 2022; Zhao et al., 2023; Yu et al., 2023; Wang et al., 2024). This common practice limits the benchmark’s ability to assess generalisation in two key ways: First, agents may overfit to specific spatial configurations, interaction bottlenecks, or object placements seen during training, without acquiring coordination strategies transferable to novel environments. Second, agents may implicitly adapt to their training partners’ behaviour patterns on known layouts, but this does not test their ability to infer intent or adapt dynamically to the behaviour of unknown partners. Crucially, these two challenges are closely linked: coordination strategies often depend on the structure of the environment. For example, in layouts with narrow passages or asymmetrical tasks, agents need carefully coordinate their movements or take on complementary roles. Testing how well agents generalise to new partners in just one environment overlooks this connection. Similarly, evaluating agents only in new environments with familiar partners misses the challenge of adapting to different behaviours. Real-world teamwork requires agents to generalise jointly over both axes.
@@
Figure 1: In the Overcooked Generalisation Challenge (OGC), during training, agents can access a generator that outputs new training environments. During evaluation, agents are presented with a novel environment and an unknown partner to cooperate with.
To address these limitations, we introduce the Overcooked Generalisation Challenge (OGC) – a novel cooperation benchmark based on Overcooked-AI to evaluate RL agents’ ability to collaborate with unknown partners and in novel environments (see Figure 1). In contrast to prior work that augments a fixed set of test environments for training Jha et al. (2025), we introduce an unsupervised environment design (UED) approach (Dennis et al., 2020), situated within the broader Dual Curriculum Design (DCD) framework, to procedurally generate a large set of diverse training layouts. DCD refers to a framework in which a generator (designer) and a level selector (curator) co-evolve to construct a curriculum tailored to the agent’s learning progress (Jiang et al., 2021a). This enables us to evaluate generalisation in a more challenging and realistic way, where agents encounter entirely novel environments without prior exposure. As such, the OGC is the first benchmark to combine UED with multi-agent zero-shot cooperation, thus bridging two previously separate lines of research.
We evaluate trained agents on a suite of human-authored test layouts across three coordination settings: self-play, cross-play, and ad-hoc teamwork. We show that OGC presents a significant challenge for current UED algorithms and scalable neural architectures. Among the evaluated methods, only PAIRED (Dennis et al., 2020) combined with a Soft Mixture-of-Experts (SoftMoE) policy (Obando-Ceron et al., 2024) achieves partial generalisation. Our results reveal a key limitation: current dual curriculum design (DCD) methods with hand-crafted or weak procedural designers fail to generate sufficiently diverse and structured training levels, and thus struggle in complex design spaces like Overcooked. By contrast, methods with learned environment generators, such as PAIRED, show better adaptability. These findings highlight the need for joint-curriculum methods that combine effective partner generalisation with adaptive curriculum generation. Our contributions are as follows:
@@
    We introduce the Overcooked Generalisation Challenge (OGC), the first open benchmark that jointly evaluates agents on environment and partner generalisation in cooperative multi-agent settings.
@@
    We release OvercookedUED – a JAX-accelerated, open-source extension of Overcooked-AI that supports unsupervised environment design (UED) via integration with state-of-the-art dual curriculum design (DCD) algorithms in minimax (Jiang et al., 2023) and JaxMARL (Rutherford et al., 2024b).
@@
    Through extensive experiments, we demonstrate that current DCD algorithms and scalable neural architectures – including recent state-of-the-art models – fail to generalise effectively across environments and partners, thus establishing OGC as a challenging new testbed for multi-agent cooperation.
@@ 2 Related Work
2.1 Partner Generalisation
Generalisation to novel partners has been studied under the ad-hoc teamwork (Stone et al., 2010) and zero-shot coordination (Hu et al., 2020) paradigms, both motivated by improving human-AI cooperation. A prominent benchmark in this space is Overcooked-AI (Carroll et al., 2019), where agents must jointly prepare and serve dishes. Many recent works use this environment to evaluate ad-hoc coordination capabilities (Strouse et al., 2021; Li et al., 2023b; Yan et al., 2023a; Liu et al., 2024; Tan et al., 2024).
In this setting, self-play often fails to produce agents that generalise to novel partners (Carroll et al., 2019). Consequently, researchers have turned to population-based methods that train diverse partner policies and learn best-response strategies in fixed environments (Zhao et al., 2023; Yu et al., 2023; Wang et al., 2024). However, these methods scale poorly, as training cost increases linearly with population size per environment Yan et al. (2023b).
In contrast, our setting trains agents across a large distribution of procedurally generated environments. Cooperation is evaluated on human-authored levels not seen during training – making population-based approaches infeasible and calling for learning strategies that operate effectively across novel partners and tasks.
2.2 Environment Generalisation
@@
RL agents fail to generalise to new environments out-of-the-box (Zhang et al., 2018a) and instead require sufficiently diverse training levels to generalise well (Zhang et al., 2018b; Cobbe et al., 2019; 2020). One established approach to generate diverse training data is domain randomisation (DR; Jakobi, 1997). However, DR may produce uninformative samples (Khirodkar et al., 2018), which hinder learning (Dennis et al., 2020).
To improve sample quality, unsupervised environment design (UED) (Dennis et al., 2020) adaptively generates levels that match an agent’s current capabilities. Prominent UED algorithms include PAIRED (Dennis et al., 2020), Prioritized Level Replay (PLR) (Jiang et al., 2021b), and ACCEL (Parker-Holder et al., 2022). These methods fall under the broader Dual Curriculum Design (DCD) framework (Jiang et al., 2021a), in which a generator and curator co-evolve to construct an adaptive training curriculum. While the development of DCD methods has been steady, they have mostly been explored in simple single-agent environments, e.g. in mazes (Dennis et al., 2020; Jiang et al., 2021a; Parker-Holder et al., 2022; Jiang et al., 2023; Li et al., 2023a; Beukman et al., 2024), bipedal walker (Wang et al., 2019; 2020; Parker-Holder et al., 2022) or car racing environments (Jiang et al., 2021a).
Multi-agent UED, in contrast, remains largely underexplored. Existing works are either closed source (Team et al., 2021; Bauer et al., 2023), do not address a (fully-)cooperative setting (Suarez et al., 2021; 2023; Samvelyan et al., 2023) or only feature multi-agent path-finding with no agent interaction (Rutherford et al., 2024a). Moreover, the underlying design spaces are shallow – often involving only walls, agents, sparse control points or pregenerated levels without design control (Dennis et al., 2020; Parker-Holder et al., 2022; Samvelyan et al., 2023; Nikulin et al., 2023).
In contrast, OGC introduces a fully cooperative multi-agent UED environment with rich object interactions (e.g., pots, onions, plates) and complex spatial dependencies (see Figure 3). The difficulty of each task critically depends on object placement, making level design substantially more challenging. The OGC thus contributes the first open-source cooperative multi-agent UED environment in which agents are exposed to novel partners during evaluation.
2.3 Combining Partner and Environment Generalisation
While many benchmarks focus on either environment or partner generalisation (Lowe et al., 2017; Foerster et al., 2018; Hu et al., 2020), few evaluate both simultaneously. A recent Overcooked study explored cross-environment cooperation (Jha et al., 2025), but their approach relied on training agents across augmented variations of known test levels, therefore implicitly assuming access to the test distribution and significantly constraining the scope of generalisation.
In contrast, the OGC poses a strictly harder challenge: agents must learn to cooperate in entirely novel environments and with unseen partners, with no prior exposure to evaluation layouts or their structure. Training is conducted solely via procedurally generated levels using UED, without handcrafted augmentations or test-level tuning. The setting of (Jha et al., 2025) can be seen as addressing a reduced version of OGC where one assumes access to the testing layouts.
This decoupled setting reveals that current DCD algorithms struggle in high-complexity cooperative domains, and motivates a new class of approaches – UED-ZSC methods – that jointly tackle unsupervised environment design and zero-shot coordination. We propose OGC as both a benchmark and a testbed to support this emerging line of research, encouraging future methods that generalise across partners and environments in realistic, open-ended tasks.
@@ 3 Preliminaries
We formalise our cooperative multi-agent UED setting as a decentralised under-specified partially observable Markov decision process (Dec-UPOMDP) with shared rewards. A Dec-UPOMDP is defined as $\mathcal{M}=\langle\mathcal{N},A,\Omega,\Theta,\mathcal{S}^{\mathcal{M}},\mathcal{T}^{\mathcal{M}},O^{\mathcal{M}},\mathcal{R}^{\mathcal{M}},\gamma\rangle$ in which $\mathcal{N}$ is the set of agents with cardinality $n$, $\Omega$ is a set of observations, and $\mathcal{S}^{\mathcal{M}}$ is the set of true states in the environment. Partial observations $o^{i}\in\Omega$ are obtained by agent $i\in\mathcal{N}$ using the observation function $O:\mathcal{S}\times\mathcal{N}\rightarrow\Omega$. Following Jiang et al. (2021a), a level $\mathcal{M}_{\theta}$ is defined as a fully-specified environment given some parameters $\theta\in\Theta$. In it, agents each pick an action $a_{i}\in A$ simultaneously to produce a joint action $\boldsymbol{a}=(a_{1},\dots,a_{n})$ and observe a shared immediate reward $R(s,\boldsymbol{a})$. Then, the environment transitions to the next state according to a transition function $\mathcal{T}:\mathcal{S}\times\mathcal{A}^{1}\times...\times\mathcal{A}^{n}\times\Theta\rightarrow\Delta(\mathcal{S})$ where $\Delta(\mathcal{S})$ refers to the space of distributions over $\mathcal{S}$. $\gamma\in[0,1)$ specifies the discount factor. Agents learn a policy $\pi$. The joint policy $\boldsymbol{\pi}$ together with the discounted return $R_{t}=\sum^{\infty}_{i=0}\gamma^{i}r_{t+1}$ induce a joint action value function $Q^{\boldsymbol{\pi}}=\mathbb{E}_{s_{t+1:\infty},\boldsymbol{a}_{t+1:\infty}}[R_{t}|s_{t},\boldsymbol{a}_{t}]$. Our formulation extends the Dec-POMDP framework (Oliehoek & Amato, 2016; Wu et al., 2021) by introducing $\Theta$ as a set of free environment parameters – making the model suitable for unsupervised environment design. This follows previous work (Dennis et al., 2020; Jiang et al., 2021a; Samvelyan et al., 2023), but differs from Samvelyan et al. (2023) in assuming shared rewards and a cooperative, rather than general-sum, structure. Within our Dec-UPOMDP, we perform UED to train a policy over a distribution of fully specified environments that enable optimal learning. This is facilitated by obtaining an environment policy $\Lambda$ (Dennis et al., 2020) that specifies a sequence of environment parameters $\Theta^{T}$ for the given policy that is to be trained. How $\Lambda$ is obtained depends on the DCD method. In OvercookedUED, $\Theta$ represents the possible positions of walls, pots, serving spots, agent starting locations, and onion and bowl piles adjusted by $\Lambda$ throughout training.